homework help true or false question.

5. If f is differentiable on R, f(1) = 8, and f(15) = 10, then there is a number c for which f(c) = 7 TRUE OR FALSE
Answer #1

on (a,b). if f is differentiable f ů continuous [a,b] from mean value theorem. where ce(a,b) fcc) = f(b) - fca) b-a b=15 f1bon (a,b). if f is differentiable f ů continuous [a,b] from mean value theorem. where ce(a,b) fcc) = f(b) - fca) b-a b=15 f1b